OpenStudy (anonymous):

1/y= 1/p+1/q solve for y.

OpenStudy (anonymous):

Is there something exactly that you don't understand...?

OpenStudy (anonymous):

how to solve for y?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Bring both fractions on the right hand side (RHS) together by finding the common denominator and adding: \[\frac{1}{y}=\frac{p+q}{pq}\] Divide each side by one \[1\div\frac{1}{y}=\frac{p+q}{pq}\div1\] This results in each fraction being flipped and thus: \[y=\frac{pq}{p+q}\]